Commissioning Notes 20080414 UT

Observers: J. Hill, R. Green, R. Zhelem, I. Foppiani in the control room and a flock of programmers in Tucson
Telescope Operator: D. Huerta

Summary

We spent the first 2/3 of the night fighting a combination of hardware problems in HBS, and some problems with the new and old version of PCS. We eventually conquerred the HBS by doing a power-off cold boot of the PLC. There is still a problem that HBS is getting insufficient flow of oil through the heat exchangers. The PCS issues got straightened out with a new build. Michele fixed the display problems on PCS that were irritated by offset commands. This new PCS does offsets/dithers in Absolute and not Relative coordinates. We measured the rotator center on IRTC by pointing at the globular cluster M3. We did only a small amount of pointing and collimation work tonight. The initial collimation model is decent. The intial pointing model is better than the one we had last night, but we still have to spiral search to find stars. We've noticed the peculiar fact that we alway find the star about 1 arcminute from where we expect, but in different directions. Seeing was 1 arcsec or better all night long. We got images in H that were 0.6 arcsec FWHM, but they are completely dominated by astigmatism and coma, so the H seeing must have been 0.4 arcsec FWHM. We spent the latter part of the night testing the optical properties of the IRTC. We even took speckle movies of a binary star in the 10 mas/pixel fine scale. Focussing motion with the hexapod was very smooth -- none of the 2 arcsec jumps that the SX primary has.

Misc. Tests in the late afternoon

Running a new version of mcstemp with thread hung messages

Running a new version of irs with the ClearStar issue solved

Running a new version of LBTtools with wavelength of 1.4 microns in pointto

Nighttime Ops

Open dome to another lovely night. Some scattered cirrus.

Mount encoders reset on the first try tonight!

Preset to M67

Software guys are experimenting with a new version of OSS, but we roll back to the old version for a while.

02:42 Copied new collimation model file to /usr/commonsw/Configuration/PSF/SXPMIRTCFCollimation.dat on lbtmu104. Remember that it is only on lbtmu104 and in /home/lbcobs/Collimation/

# First collimation and pointing data with SX Front Bent Gregorian IRTC
# Hill, Green, Zhelem and Foppiani
# Temperature was +5 degC
#irtc2_20080413.002505.fits 0.4427857 -2.03432 1.1895513 24.21538 -29.97316 -0.01810467 39.0
#irtc2_20080413.014205.fits -0.3538426 -1.99989 -0.1069624 18.86937 -29.97085 -0.01392317 74.1
#irtc2_20080413.022502.fits -0.7101959 -1.828912 -0.4062092 17.68564 -29.89798 0.001143336 82.0
#irtc2_20080413.024923.fits -0.5102229 -1.772641 -0.4215685 17.32344 -29.91453 0.03437164 84.1
#irtc2_20080413.042747.fits -0.1620223 -2.454743 0.277193 19.34836 -30.01535 0.02929986 72.6
#irtc2_20080413.045257.fits 0.481838 -2.412749 1.385125 24.46381 -29.99549 0.01050852 38.1
#irtc2_20080413.045729.fits 0.4813082 -2.413064 1.385126 24.38936 -29.94351 0.05277823 38.1

~02:50 Telescope stops tracking because of HBS problems until 03:02

02:58 Michelle installing new version of PCS until 03:03, but it had RPC problems.

03:28 Reverted to last night's PCS, and go back to M67,

03:30 HBS stops again. (too bad it is anti-correlated with PCS) David is rebooting the WinCC machine. That worked until 04:08, when it stopped again.

In the meantime, the rolled back PCS wouldn't work, and this has degenerated into a full-blown TCS debugging section.

05:40 It took a power off cold boot of HBS PLC to get the alarms to stop, plus the usual forces to deal with E31 and E3 comp. HBS at least for a while is running OK. Oil Tank temperature is 17 degC, and we have the glycol valve open manually to 65% trying to get things to cool off.

~05:50 Issuing a preset from the preset GUI make PCS die (either Michelle's normal version or the standard build that we were using last night).

Now we have a new pointing model from Dave T. /usr/commonsw/Configuration/PCS/irtc.20080414.ptmod (that wasn't Dave's name which was irtc.mask5.ptmod). This is a 4-star fit with a 5th suspect star deleted.

Finally tracking near globular cluster M3, but lots of strange things......

Mirror or telescope or something seems to have a residual drift after a repointing. Maybe because CSQ isn't waiting for rotator. Or maybe our minimalist pointing model isn't so good?

IRTC.offset commands don't seem reversible. (see below)

Focus and coma are mushy from one re-collimation to the next.

07:34 sending offset 0 0 ROToff=30 moves rotator from 111 to 141, but ROToff=-30 moves back from 141 to 80.
irc OffsetPointing -30. 0. 0. RADEC_FOCAL MOUNT true REL left

Preset to M13 with irtc.20080413 16 41 41.44 +36 27 36.9

Discovered that in tonight's PCS offsets are absolute.

Measuring the IRTC rotator center

08:32 Pair of M13 images with rotator offsets of 0 and 180, followed by a short movie of rotating back to 0

08:34 Back to today's pointing model irtc.20080414.ptmod (mask5). This pointing model has significantly less drift than the one from yesterday.

08:50 At -70 -55 in M13, record movie of images apparently oscillating in 1 axis, perhaps it is the atmosphere combined with some astigmatism.

08:54 Take still images of 2, 1, and 3 sec in this field (wide mode). Images are ~0.6 arcsec FWHM but completely dominated by astigmatism and coma, so the seeing must be about 0.4 arcsec at H.

09:03 IRTC GUI crashes while recording some long integrations. Restart it and reset temperature to 245K.

09:10 focus jump while tracking M13, while readjusting sometimes 10 microns of M1 Z-motion makes 2 arcsec of lateral shift of the F/15 image.

09:14 takepic 1 sec image in M13 after refocus

09:15 pointto WT10_228, Search with spiral found at -53,63

*09:30 Image for collimation and pointing at EL=62.4, but maybe the GetImage failed?

IRTC camera optical tests, and camera operation training

Preset to double star 15:50:53.3 19:10:48.9 at EL=75.3, spiral search found at dRA,dDEC -12,64 (again the image is ~1 arcminute off)

10:01 0.5 sec image on double star, M1 global offsets of X=-0.08 Y-0.09 Z=0.12 RX=0 RY=0 (so the new collimation model works here)

Italo is zooming to the medium 20 mas/pixel scale, telescope is drifting West in RA at this position.

10:23 Preset to brighter double star: 15:42:44.5 +26:17:44.3, spiral search found at 27,52

M1 global offsets X=-0.08, Y=0.1, Z=0.13, RX=0, RY=0

10:38 Italo took some speckle movies of this pair

10:44 Zoom to the fine scale at 10 mas/pixel. Seeing monitor says 1.0 arcsec during these movies.

10:47 Test without the pupil stop, IRTC GUI crashed

Focussing at the 10 mas scale with hexpod shows no image position jitter above the atmospheric motion. (unlike M1 which has 2 arcsec jitter) Sweet. Tracking here at EL=69.3 is quite stable, drift of maybe 1 arcsec in 10 minutes.

11:03 Telescope stops tracking because of lost polynomials. (Michele says one instance of this is due to timing between PCS and MCSPU.)

11:04 Stephen gives us new PCS and new OSS, but new OSS isn't speaking to hexapod, so back out the OSS.

11:08 Telescope stops tracking again because of lost polynomials.

11:24 OSS is reconnected to hexapod and we are back on target, too bad we can't see the target at the old offset.

Jose gave us a new version of irs with longer socket timeouts. (before about 5% of camera interactions were failing) Initial operation was OK, but we need to test it more tomorrow night.

11:30 Start another spiral search. Michelle's new PCS is now handling the offsets correct and is reporting sensible achieved coordinates in the DD. Found the star near 22,90. WHY is it different than it was an hour ago by 40 arcsec?

11:43 Preset again with PA=180, and star is at the same offset (22,90).

11:47 pointto WT10_249, spiral search --- We saw it go by twice, but couldn't backtrack to it

12:03 pointo WT10_249 again, found at -15,70

12:10 takepic 0.5

Correlated motion test to confirm the signs of M1 and M2 motion in absolute space.

Move M1 X -0.5 this needs M2 X +0.5 to make no coma

Move M1 Z -0.5 M2 Z -0.5 cancels that defocus Means -Z on M2 moves toward primary.

Move M1 Y +0.5 this needs M2 Y -0.5 to make no coma

Conclusion: Both are right hand systems, but M2 coordinates need to be rotated 180 deg around Z-axis to match the primary coordinates. (They were rotated the following day.)

Ross has measured the relation of the angles to the X,Y axes with the alignment telescope.

12:22 Close dome, and park telescope at horizon

David has gone to power cycle the MAT.

-- JohnHill - 14 Apr 2008
Topic revision: r7 - 19 Jan 2009, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback